HIGH FRECQUENCY NiZn FERRITE PREPARED BY SELF-COMBUSTION METHOD WITH NONMAGNEIC ADDITIVES

N.Rezlescu, P.D.Popa and E.Rezlescu

Institute of Technical Physics, Bd.D.Mangeron 47, 6600 Iasi, Romania

It is known the application of NiZn ferrite for making transducer heads for magnetic recording as a substitute for laminated metallic alloys, in order to improve wear resistance and to reduce the losses. The features of this ferrite are the high mechanical strengh, high hardness, high resistivity, chemical stability and reasonable cost.

The main problem in preparing these polycrystalinne materials is the porosity. This should be less than 1% [1]. To overcome the problem of the density and to optimise the numerous properties of interest such as coercivity, permeability, resistivity, we are investigated the influence of the various nonmagnetic additives as Na_{2}O, CaO, Sb_{2}O_{3}, ZrO_{2}, K_{2}O and Li_{2}O. Using self combustion procedure and materials of high purity, we have succeded in preparing of polycrystalline ferrite, Ni_{0.36}Zn_{0.64}Fe_{2}O_{4} with a reduced porosity of about 1.9% . The increase in the density with the increasing additive amount was explained by additives segregation on the grain boundaries. Regarding the others properties, we have found that the CaO and Na_{2}O increase remarkably the electrial resistivity while Li_{2}O increases the permeability. From the additives investigated, the triple compositions Na_{2}O- CaO-Li_{2}O and Na_{2}O-CaO-ZrO_{2}led to the best results regarding the magnetic and mechanical characteristics of NiZn ferrite acceptable for magnetic head.

[1] E.Hirota et all., Proc.Intern.Conf.Ferrites, Kyoto, Japan, 1980, p.667.

Presentation: poster

Abstract categorie: 3 – Transition Metals, Surfaces and Multilayers Corresponding authors: Nicolae REZLESCU Institute of Technical Physics B-dul D.Mangeron 47 Iasi 6600 ROMANIA Tel: 40-32-130680 Fax: 40-32-231132 E-mail: reznic@phys-iasi.ro